# Symmetry-Informed Governing Equation Discovery

### v1

@@ Abstract
Despite the advancements in learning governing differential equations from observations of dynamical systems, data-driven methods are often unaware of fundamental physical laws, such as frame invariance. As a result, these algorithms may search an unnecessarily large space and discover equations that are less accurate or overly complex. In this paper, we propose to leverage symmetry in automated equation discovery to compress the equation search space and improve the accuracy and simplicity of the learned equations. Specifically, we derive equivariance constraints from the time-independent symmetries of ODEs. Depending on the types of symmetries, we develop a pipeline for incorporating symmetry constraints into various equation discovery algorithms, including sparse regression and genetic programming. In experiments across a diverse range of dynamical systems, our approach demonstrates better robustness against noise and recovers governing equations with significantly higher probability than baselines without symmetry.
@@ 1 Introduction
An important aspect of equation discovery is to ensure that the discovered equations adhere to physical principles, such as conservation laws and symmetries. Existing methods that fail to consider these principles may end up searching in an unnecessarily large space of equations and discover equations that do not conform to fundamental physical laws.
@@
  • Overall, we establish a pipeline to use Lie point symmetries of ODEs to constrain the equation learning problem.
  • We theoretically derive the criterion for symmetry of time-independent ODEs in terms of equivariance of the associated flow map.
  • From this criterion, we solve the constraint explicitly for compressing the equation search space in sparse regression and promoting parsimony in the learned equations.
  • Alternatively, we promote symmetry by a symmetry regularization term when the symmetry constraint cannot be explicitly solved.
  • In experiments across many dynamical systems with substantial noise, our symmetry-informed approach achieves higher success rates in recovering the governing equations.
@@ 2.2 Symmetry Prior in Machine Learning
Symmetry plays a crucial role in machine learning. There has been a vast body of work on equivariant neural networks (Cohen & Welling, 2016; Weiler & Cesa, 2019; Finzi et al., 2020, 2021; Bronstein et al., 2021; Wang et al., 2021). These network architectures enforce various symmetries in different data types. Symmetry can also be exploited through data augmentation (Benton et al., 2020) and canonicalization (Kaba et al., 2023). There are already some works that exploit symmetries for learning in dynamical systems governed by differential equations through data augmentation (Brandstetter et al., 2022), symmetry loss (Akhound-Sadegh et al., 2023) and contrastive learning (Mialon et al., 2023). However, works using symmetry in recovering underlying equations from data are scarce, and existing examples only consider symmetries specific to a particular system, e.g. reflections and permutations in proper orthogonal decomposition (POD) coefficients (Guan et al., 2021), rotations and translations in space (Ridderbusch et al., 2021; Baddoo et al., 2021), etc. In this work, we provide solutions for dealing with general matrix Lie group symmetries.
While symmetry has proved an important inductive bias, for the ODEs considered in this work, the underlying symmetries are often not known a priori. Some recent works have aimed to automatically discover symmetries from data (Liu & Tegmark, 2022; Yang et al., 2023b, a; Otto et al., 2023). We use the adversarial framework in Yang et al. (2023a) to learn unknown symmetries from data and discover the equations subsequently using the learned symmetry.
@@ 3 Symmetry of ODEs
Our goal is to discover the governing equations of a dynamical system (1) in symbolic form. Our dataset consists of observed trajectories of the system, $\{\bm{x}_{0:T}^{(i)}\}_{i=1}^{N}$, where $\bm{x}_{t}^{(i)}\in X\subseteq\mathbb{R}^{d}$ denotes the observation at timestep $t$ in the $i$th trajectory. We assume the observations are collected at a regular time interval $\Delta t$. In our methodology discussion, we omit the trajectory index $i$ and consider one trajectory, $\bm{x}_{0:T}$ for simplicity. To leverage symmetry for better equation discovery, we first define symmetries in ODEs and provide the theoretical foundation for our methodology. In particular, we show that a time-independent symmetry of the ODE is equivalent to the equivariance of its flow map. This equivariance property can then be conveniently exploited to enforce symmetry constraints in the equation discovery process.
@@ 3.1 Lie Point Symmetry and Flow Map Equivariance
We refer the readers to Appendix A for background on Lie groups, group actions, and their applications on differential equations. In there, We also discuss the general time-dependent symmetries for (non-autonomous) ODEs, which provides the basis for the potential generalization of our method. For now, we consider only time-independent symmetries in Definition 3.1 and demonstrate how to use these symmetries to derive equivariance constraints on equations.
@@ 3.2 The Infinitesimal Formulae
Let $G$ be a time-independent symmetry group of the ODE (1). Let $v\in\mathfrak{g}$ be an element in the Lie algebra $\mathfrak{g}$ of the group $G$. Consider the flow map $\bm{f}$ defined in (2) for a fixed time interval. Denote $J_{\bm{f}}$ and $J_{g}$ as the Jacobian of $\bm{f}(\bm{x})$ and $g\cdot\bm{x}$ w.r.t $\bm{x}$. For all $\bm{x}\in X$ and $g=\exp(\epsilon v)\in G$,
@@
To interpret Theorem 3.3, (3.3) follows directly from Proposition 3.2, and the other equations are the infinitesimal equivalent of this equivariance condition. For example, the Jacobian-vector product in (3.3) reveals how much $\bm{f}$ changes when the input is infinitesimally transformed by $v$, and $v(\bm{f}(\bm{x}))$ indicates the amount of change caused by the infinitesimal action on the output of $\bm{f}$.  [^1] Note that $v(\bm{f}(\bm{x}))$ means evaluating the vector field at $\bm{f}(\bm{x})\in X$. It is different from $v(\bm{f})(\bm{x})$, which is the Lie derivative of $\bm{f}$ at the point $\bm{x}$. The difference between these two terms vanishes if the function is equivariant. On the other hand, as both $\bm{h}$ and $v$ can be viewed as vector fields on $X$, their roles are interchangeable in the equivariance formula. Therefore, we can also express the equivariance in terms of the flow $\bm{h}$ and a finite group element $g=\exp(\epsilon v)$. Finally, we can also consider the fully infinitesimal formula (3.3), which indicates that $\bm{h}$ and $v$ commute as vector fields. The proof of Theorem 3.3 may be found in Section B.1.
@@ 4 Symmetry-Informed Governing Equation Discovery
Next, we discuss how to leverage the knowledge of symmetry in equation discovery. Depending on the nature of different symmetries, we develop a pipeline for incorporating symmetries into various equation discovery algorithms, shown in Figure 1. We introduce the components in the following subsections.
Figure 1: Pipeline for incorporating symmetries into equation discovery via solving linear symmetry constraint (Section 4.1), regularization (Section 4.2) and symmetry discovery (Section 4.3).
@@ 4.2 Equiv-r: Regularization for General Symmetry
We use a relative loss, because the scale of both terms in the numerator and their difference is subject to the specific $\bm{f}_{\tau}$. In the extreme case when $\bm{h}=\bm{0}$, both terms are zero because $\bm{f}_{\tau}(\bm{x})$ is constant. As we are optimizing $\bm{h}$ under this objective, we compute the relative error to eliminate the influence of the level of variations in the function itself on the symmetry error.
@@
The use of symmetry as a regularization term is a general approach that can be applied to various equation discovery algorithms, e.g. sparse regression and genetic programming. Similar to the works that utilize the variational form of ODEs (Messenger & Bortz, 2021; Qian et al., 2022), our symmetry regularization does not require estimating the time derivative, so it is more robust to noise. We name this approach as EquivSINDy-r (or EquivGP-r for genetic programming-based discovery algorithms), r standing for regularization.
@@ 5 Experiments
For each ODE system, we sample $N$ random initial conditions $\bm{x}_{0}$ from a uniform distribution on a specified subset of $X\subseteq\mathbb{R}^{d}$. Starting from each initial condition, we integrate the ODE using the RK4 method and sample with a regular step size $\Delta t$ to get the discrete trajectory $\bm{x}_{0:T}$. We use different $N$, $T$ and $\Delta t$ for each system, as specified in Appendix D.
@@
We run each algorithm multiple times and report the success probability of discovering the correct function form. Specifically, assume the true equation is expanded as $\sum\theta_{f_{i}}f_{i}(\bm{x})$, where $\theta_{f_{i}}$ is a nonzero constant parameter and $f_{i}$ is a function of $\bm{x}$. Also, we expand the discovered equation as $\sum\hat{\theta}_{f_{i}}\hat{f}_{i}(\bm{x})$, where $\hat{\theta}_{f_{i}}\not=0$. The discovery is considered successful if all the terms match, i.e. $\{f_{i}\}=\{\hat{f}_{i}\}$. The probability is computed as the proportion of successful runs. As each dynamical system consists of multiple equations (one for each dimension), we evaluate the success probability of discovering each individual equation (e.g. "Eq. 1", "Eq. 2" columns in Table 1) and the joint success probability of simultaneously discovering all the equations (e.g. "All" column in Table 1). We argue that this is the most important metric for evaluating the performance of an equation discovery algorithm, since an accurate equation form reveals the key variables and their interactions, and thus the underlying structure and relationships governing the system.
@@ 5.1 Equivariant SINDy for Linear Symmetries
In this section, we assume the knowledge of some common linear symmetries and build equivariant models. We consider the following two synthetic systems.
@@
This is a two-dimensional system with rotation symmetry, characterized by the infinitesimal generator $v=x_{2}\partial_{1}-x_{1}\partial_{2}$. It is governed by (5.1).
@@
Table 1 displays the success probability of different methods. With the reduced weight space, our equivariant model almost always recovers the correct equation forms. Additional results, including the parameter estimation error and the prediction error with discovered equations, are deferred to Section C.1. Overall, these results fully demonstrate the advantage of applying symmetry to compress the parameter space.
@@ 5.3 Symmetry Regularization
Table 2: Equation discovery statistics on Lotka-Volterra system (15) at noise level $\sigma_{R}=99\%$ and Sel’Kov oscillator (16) at $\sigma_{R}=20\%$. The success probability is computed from 50 runs for sparse regression-based algorithms and 10 runs for genetic programming-based algorithms. Full results are available in Table 4.
System Method Success prob.
Eq. 1 Eq. 2 All
L-V (15) SINDy 0.40 0.64 0.24
WSINDy 0.18 0.22 0.06
EquivSINDy-r 0.54 0.58 0.36
GP 1.0 0.0 0.0
D-CODE 0.0 0.0 0.0
EquivGP-r 1.0 0.8 0.8
Sel’Kov (16) SINDy 0.30 0.56 0.14
WSINDy 0.06 0.14 0.04
EquivSINDy-r 0.40 0.70 0.28
GP 0.0 0.0 0.0
D-CODE 0.0 0.4 0.0
EquivGP-r 0.1 0.6 0.1
@@
As we assume no knowledge of the symmetries here, we first run LaLiGAN (Yang et al., 2023a) to discover symmetry from data for each system. We then extract the group action and the induced infinitesimal action from LaLiGAN and use them to compute symmetry regularization. Table 2 shows the results of different discovery algorithms with and without symmetry. Comparison is made within each class of methods, i.e. sparse regression and genetic programming. For genetic programming, we use an alternate form of symmetry regularization (36), which will be discussed in Section C.2. Informed by symmetries, equation discovery algorithms can successfully discover equations that they are unlikely to discover otherwise. In Section C.2, we provide full results to show that our method can generally achieve lower parameter estimation error and prediction error. We also compare the different regularization options based on Theorem 3.3.
@@ Acknowledgement
This work was supported in part by the U.S. Army Research Office under Army-ECASE award W911NF-23-1-0231, the U.S. Department Of Energy, Office of Science, IARPA HAYSTAC Program, CDC-RFA-FT-23-0069, DARPA AIE FoundSci, NSF Grants #2205093, #2146343, and #2134274.

### v2

@@ Abstract
Despite the advancements in learning governing differential equations from observations of dynamical systems, data-driven methods are often unaware of fundamental physical laws, such as frame invariance. As a result, these algorithms may search an unnecessarily large space and discover less accurate or overly complex equations. In this paper, we propose to leverage symmetry in automated equation discovery to compress the equation search space and improve the accuracy and simplicity of the learned equations. Specifically, we derive equivariance constraints from the time-independent symmetries of ODEs. Depending on the types of symmetries, we develop a pipeline for incorporating symmetry constraints into various equation discovery algorithms, including sparse regression and genetic programming. In experiments across diverse dynamical systems, our approach demonstrates better robustness against noise and recovers governing equations with significantly higher probability than baselines without symmetry. Our codebase is available at https://github.com/Rose-STL-Lab/symmetry-ode-discovery.
@@ 1 Introduction
An important aspect of equation discovery is the principle of parsimony (Sober, 1981)the scientific principle that the most acceptable explanation of a phenomenon is the simplest. It promotes the scientific goal of identifying a model that offers robust predictive capabilities while maintaining simplicity. Existing methods that fail to consider these principles may end up searching in an unnecessarily large space of equations and discover equations that are overly complex or do not conform to fundamental physical laws.
@@
  • We establish a holistic pipeline to use Lie point symmetries of ODEs to improve the accuracy and robustness of equation discovery algorithms.
  • We derive the theoretical criterion for symmetry of time-independent ODEs in terms of equivariance of the associated flow map.
  • From this criterion, we solve the linear symmetry constraint explicitly for compressing the equation search space in sparse regression and promoting parsimony in the learned equations.
  • For general symmetry, we promote symmetry by a symmetry regularization term when the symmetry constraint cannot be explicitly solved.
  • In experiments across many dynamical systems with substantial noise, our symmetry-informed approach achieves higher success rates in recovering the governing equations and better robustness against noise.
@@ 2.2 Symmetry Prior in Machine Learning
Symmetry plays a crucial role in machine learning. There has been a vast body of work on equivariant neural networks (Cohen & Welling, 2016; Weiler & Cesa, 2019; Finzi et al., 2020, 2021; Bronstein et al., 2021; Wang et al., 2021). These network architectures enforce various symmetries in different data types. Symmetry can also be exploited through data augmentation (Benton et al., 2020) and canonicalization (Kaba et al., 2023). There are already some works that exploit symmetries for learning in dynamical systems governed by differential equations through data augmentation (Brandstetter et al., 2022), symmetry loss (Huh et al., 2020; Akhound-Sadegh et al., 2023) and contrastive learning (Mialon et al., 2023). However, works using symmetry in recovering underlying equations from data are scarce, and existing examples only consider symmetries specific to a particular system, e.g. reflections and permutations in proper orthogonal decomposition (POD) coefficients (Guan et al., 2021), rotations and translations in space (Ridderbusch et al., 2021; Baddoo et al., 2021), etc. In this work, we provide solutions for dealing with general matrix Lie group symmetries.
While symmetry has proved an important inductive bias, for the ODEs considered in this work, the underlying symmetries are often not known a priori. Some recent works have aimed to automatically discover symmetries from data (Liu & Tegmark, 2022; Yang et al., 2023b, a; Otto et al., 2023; van der Ouderaa et al., 2024). When there is no available prior knowledge about symmetry, We use the adversarial framework in Yang et al. (2023a) to learn unknown symmetries from data and discover the equations subsequently using the learned symmetry.
@@ 3 Symmetry of ODEs
Our goal is to discover the governing equations of a dynamical system (1) in symbolic form. Our dataset consists of observed trajectories of the system, $\{\bm{x}_{0:T}^{(i)}\}_{i=1}^{N}$, where $\bm{x}_{t}^{(i)}\in X\subseteq\mathbb{R}^{d}$ denotes the observation at timestep $t$ in the $i$th trajectory. We assume the observations are collected at a regular time interval $\Delta t$. We omit the trajectory index $i$ and consider one trajectory, $\bm{x}_{0:T}$ for simplicity. To leverage symmetry for better equation discovery, we first define symmetries in ODEs and provide the theoretical foundation for our methodology. In particular, we show that a time-independent symmetry of the ODE is equivalent to the equivariance of its flow map. This equivariance property can then be conveniently exploited to enforce symmetry constraints in the equation discovery process.
@@ 3.1 Lie Point Symmetry and Flow Map Equivariance
We refer the readers to Appendix A for background on Lie groups, group actions, and their applications on differential equations. We consider time-independent symmetries in Definition 3.1 and demonstrate how to use these symmetries to derive equivariance constraints on equations.
@@ 3.2 The Infinitesimal Formulae
Let $G$ be a time-independent symmetry group of the ODE (1). Let $v\in\mathfrak{g}$ be an element in the Lie algebra $\mathfrak{g}$ of the group $G$. Consider the flow map $\bm{f}$ defined in (2) for a fixed time interval. Denote $J_{\bm{f}}$ and $J_{g}$ as the Jacobian of $\bm{f}(\bm{x})$ and $g\cdot\bm{x}$ w.r.t $\bm{x}$. For all $\bm{x}\in X$ and $g=\exp(\epsilon v)\in G$, the following equalities hold:
@@
To understand Theorem 3.3, (3.3) follows directly from Proposition 3.2, and the other equations are the infinitesimal equivalent of this equivariance condition. For example, the Jacobian-vector product in (3.3) reveals how much $\bm{f}$ changes when the input is infinitesimally transformed by $v$, and $v(\bm{f}(\bm{x}))$ indicates the amount of change caused by the infinitesimal action on the output of $\bm{f}$.  [^1] Note that $v(\bm{f}(\bm{x}))$ means evaluating the vector field at $\bm{f}(\bm{x})\in X$. It is different from $v(\bm{f})(\bm{x})$, which is the Lie derivative of $\bm{f}$ at the point $\bm{x}$. The difference between these two terms vanishes if the function is equivariant. On the other hand, as both $\bm{h}$ and $v$ can be viewed as vector fields on $X$, their roles are interchangeable in the equivariance formula. Therefore, we can also express the equivariance in terms of the flow $\bm{h}$ and a finite group element $g=\exp(\epsilon v)$. Finally, we can also consider the fully infinitesimal formula (3.3), which indicates that $\bm{h}$ and $v$ commute as vector fields. The proof of Theorem 3.3 may be found in Section B.1.
@@ 4 Symmetry-Informed Governing Equation Discovery
Next, we discuss how to leverage the knowledge of symmetry in equation discovery. Depending on the nature of different symmetries, we develop a holistic pipeline for incorporating symmetries into various equation discovery algorithms, shown in Figure 1. We introduce the components in the following subsections.
Figure 1: Pipeline for incorporating symmetries into equation discovery via solving linear symmetry constraint (Section 4.1), regularization (Section 4.2) and symmetry discovery (Section 4.3). Given the trajectory data from the dynamical system, we first identify its symmetry based on prior knowledge or symmetry discovery techniques. We then enforce the symmetry by solving a set of constraints when possible and otherwise promote the symmetry through regularization.
@@ 4.2 Equiv-r: Regularization for General Symmetry
We use a relative loss because the scale of both terms in the numerator and their difference is subject to the specific $\bm{f}_{\tau}$. In the extreme case when $\bm{h}=\bm{0}$, both terms are zero because $\bm{f}_{\tau}(\bm{x})$ is constant. As we are optimizing $\bm{h}$ under this objective, we compute the relative error to eliminate the influence of the level of variations in the function itself on the symmetry error.
@@
The use of symmetry regularization encourages equation discovery models to achieve lower symmetry error, instead of explicitly constraining the parameter space as in Section 4.1. However, it is a general approach that can be applied to various equation discovery algorithms, e.g. sparse regression and genetic programming. Similar to the works that utilize the variational form of ODEs (Messenger & Bortz, 2021; Qian et al., 2022), our symmetry regularization does not require estimating the time derivative, so it is more robust to noise. We name this approach as EquivSINDy-r (or EquivGP-r for genetic programming-based discovery algorithms), r standing for regularization.
@@ 4.3 Equation Discovery with Unknown Symmetry
We should note that our method has a completely different goal from LaLiGAN. Our method aims to discover equations using symmetry as an inductive bias. LaLiGAN aims to discover unknown symmetry. Only when we do not know the symmetry a priori, we use LaLiGAN as a tool to discover the symmetry first and use the discovered symmetry to regularize equation learning.
@@ 5 Experiments
For each ODE system, we sample $N$ random initial conditions $\bm{x}_{0}$ from a uniform distribution on a specified subset of $X\subseteq\mathbb{R}^{d}$. Starting from each initial condition, we integrate the ODE using the 4th-order Runge-Kutta (RK4) method and sample with a regular step size $\Delta t$ to get the discrete trajectory $\bm{x}_{0:T}$. We use different $N$, $T$ and $\Delta t$ for each system, as specified in Appendix D.
@@
We run each algorithm multiple times and report the success probability of discovering the correct function form. Specifically, assume the true equation is expanded as $\sum\theta_{f_{i}}f_{i}(\bm{x})$, where $\theta_{f_{i}}$ is a nonzero constant parameter and $f_{i}$ is a function of $\bm{x}$. Also, we expand the discovered equation as $\sum\hat{\theta}_{f_{i}}\hat{f}_{i}(\bm{x})$, where $\hat{\theta}_{f_{i}}\not=0$. The discovery is considered successful if all the terms match, i.e. $\{f_{i}\}=\{\hat{f}_{i}\}$. The probability is computed as the proportion of successful runs. As each dynamical system consists of multiple equations (one for each dimension), we evaluate the success probability of discovering each individual equation (e.g. "Eq. 1", "Eq. 2" columns in Table 1) and the joint probability of successfully discovering all the equations (e.g. "All" column in Table 1). We argue that this is the most important metric for evaluating the performance of an equation discovery algorithm, since an accurate equation form reveals the key variables and their interactions, and thus the underlying structure and relationships governing the system.
@@ 5.1 Equivariant SINDy for Linear Symmetries
In this section, we consider two dynamical systems which possess some linear symmetries. These symmetries are commonly observed across many systems and their presence can be easily detected. Therefore, we assume we know these symmetries and focus on symmetry-informed equation discovery by solving the symmetry constraints.
@@
This is a two-dimensional system with rotation symmetry, characterized by the infinitesimal generator $v=x_{2}\partial_{1}-x_{1}\partial_{2}$ (i.e. $\begin{bmatrix}0&1\\ -1&0\end{bmatrix}$ in matrix terms). It is governed by (5.1).
@@
Table 1 displays the success probability of different methods. With the reduced weight space, our equivariant model almost always recovers the correct equation forms. Full results, including the parameter estimation error and the prediction error with discovered equations, as well as the experiment on a higher-dimensional system, are deferred to Section C.1. Overall, these results fully demonstrate the advantage of applying symmetry to compress the parameter space.
@@ 5.3 Symmetry Regularization
Table 2: Equation discovery statistics on the Lotka-Volterra system (15) at noise level $\sigma_{R}=99\%$ and the glycolytic oscillator (16) at noise level $\sigma_{R}=20\%$. The RMSE is scaled by ×$\times 10^{-1}$ for (15) and ×$\times 10^{-2}$ for (16). The success probability is computed from 50 runs for sparse regression-based algorithms and 10 runs for genetic programming-based algorithms. The success probabilities of recovering individual equations (Eq. 1 & Eq. 2) and simultaneously recovering both equations (All) are reported. The RMSE (all) refers to the parameter estimation error over all runs. The RMSE (successful) refers to the parameter estimation error over successful runs, which is missing for algorithms with zero success probability.
System Method Success prob. RMSE (successful) RMSE (all)
Eq. 1 Eq. 2 All Eq. 1 Eq. 2 All Eq. 1 Eq. 2
GP 1.0 0.0 0.0 2.44 (0.89) N/A N/A 2.44 (0.89) 13.20 (3.20)
D-CODE 0.0 0.0 0.0 N/A N/A N/A 10.38 (0.11) 9.08 (1.88)
Glycolytic Oscillator (16) SINDy 0.30 0.56 0.14 0.87 (0.15) 0.32 (0.13) 0.67 (0.10) 15.86 (17.52) 12.71 (18.32)
GP 0.0 0.0 0.0 N/A N/A N/A 7.57 (5.46) 14.88 (23.50)
D-CODE 0.0 0.4 0.0 N/A 0.27 (0.11) N/A 14.37 (15.66) 8.90 (16.84)
@@
Since we assume no knowledge of the symmetries here, we first run LaLiGAN (Yang et al., 2023a) to discover symmetry from data for each system. We then extract the group action and the induced infinitesimal action from LaLiGAN and use them to compute symmetry regularization. Table 2 shows the results of different discovery algorithms with and without symmetry. Comparison is made within each class of methods, i.e. sparse regression and genetic programming. For genetic programming, we use an alternate form of symmetry regularization (37), which will be discussed in Section C.2.
Informed by symmetries through regularization terms, equation discovery algorithms have a much higher success probability of discovering the correct equation forms. Also, models with symmetry regularization generally have the lowest parameter estimation errors averaging over all random experiments.
In Section C.2, we provide additional results to show that the equations discovered by our method also achieve lower prediction error. We also compare the different regularization options based on Theorem 3.3.
@@ Acknowledgement
This work was supported in part by the U.S. Army Research Office under Army-ECASE award W911NF-23-1-0231, the U.S. Department Of Energy, Office of Science, IARPA HAYSTAC Program, CDC-RFA-FT-23-0069, DARPA AIE FoundSci, DARPA YFA, NSF Grants #2205093, #2146343, and #2134274.